# Subdiffusive concentration for the chemical distance in Bernoulli percolation

### v1

@@ 1.4. Organization and notation of this paper
  • Integer interval. Given an integer $t\geq 1$, we denote by $[t]:=\{1,2,\ldots,t\}$.
@@
  • Box and its boundary. Let $x\in{\mathbb{Z}}^{d}$ and $t>0$, we will denote by $\Lambda_{t}(x):=x+[-t,t]^{d}\cap{\mathbb{Z}}^{d}$ the box centered at $x=(x_{1},\ldots,x_{d})\in{\mathbb{Z}}^{d}$ with radius $t$. For convenience, we briefly write $\Lambda_{t}=[-t,t]^{d}$ for $\Lambda_{t}(0)$. We define the boundary of $\Lambda_{t}(x)$ as $\partial\Lambda_{t}(x):=\Lambda_{t}(x)\setminus\Lambda_{t-1}(x)$.
@@
  • ${\mathbb{Z}}^{d}$-path and set of ${\mathbb{Z}}^{d}$-paths. For any $\ell$, we say that a sequence $\gamma=(v_{0},\ldots,v_{\ell})$ is a ${\mathbb{Z}}^{d}$-path if for all $i\in[\ell],\|v_{i}-v_{i-1}\|_{1}=1$. The length of $\gamma$ is $\ell$, denoted by $|\gamma|$. For $1\leq i<j\leq\ell$, we denote by $\gamma_{v_{i},v_{j}}$ the sub-path of $\gamma$ from $v_{i}$ to $v_{j}$. In addition, if $v_{i}\neq v_{j}$ for $i\neq j$, then we say that $\gamma$ is self-avoiding. From now on, we will shortly write a path in place of a self-avoiding ${\mathbb{Z}}^{d}$-path. For any path $\gamma$, we denote ${\bf s}(\gamma),{\bf e}(\gamma)$ the starting and ending vertices of $\gamma$, respectively. Given $U\subset{\mathbb{Z}}^{d}$, let $\mathcal{P}(U)$ be the set of all paths in $U$.
@@
[…]
@@
    and we thus denote diam$(A)$ the diameter of $A$ by
@@
  • Geodesic. Given $\{x,y\}\subset U\subset{\mathbb{Z}}^{d}$, we denote by $\gamma$ a geodesic between $x$ and $y$ of ${\mathrm{D}}^{U}(x,y)$ if $\gamma$ is an open path inside $U$ such that $|\gamma|={\mathrm{D}}^{U}(x,y)$. We also denote by $\eta$ a geodesic between $x$ and $y$ of ${\mathrm{T}}^{U}(x,y)$ if $\eta$ is a path inside $U$ such that ${\mathrm{T}}(\eta)={\mathrm{T}}^{U}(x,y)$. If there are several choices for $\gamma$ or $\eta$, we choose one of them by a deterministic rule.
@@ 2. The effect of resampling
As mentioned in Introduction, the key to the concentration of the truncated passage time ${\mathrm{T}}_{n}$ is understanding the cost of resampling the edge weights along the geodesic of ${\mathrm{T}}_{n}$. To study this issue, given an edge $e$, we introduce the notion of effective radius $R_{e}$, which measures how large the change in passage time is when flipping the weight of $e$. The radius $R_{e}$ indeed guarantees an open bypass of $e$ whose length is comparable to $R_{e}$. Hence, it turns out that the cost of flipping the edge weight is given by (1.12). We then investigate the properties of the effective radii, including the local dependence and light-tailed decay distribution. These properties will play a central role in controlling the total cost of resampling edges.
@@ 2.1. Effective radius and its applications
$\mathbb{G}_{*}(u,v;A)$ ::=:= $\Big{\{}\gamma\in{\mathcal{P}}(A):{\bf s}(\gamma)=u,{\bf e}(\gamma)=v,\exists\,\pi\in\mathbb{G}(A):\gamma\setminus\pi\textrm{ is open}\Big{\}},$
$\mathbb{G}_{*}(A)$ ::=:= $\bigcup_{u,v\in A}\mathbb{G}_{*}(u,v;A),$
@@
Given $B\subset A\subset{\mathbb{Z}}^{d}$, if $\gamma\in\mathbb{G}(A)$ (or $\gamma\in\mathbb{G}_{*}(A)$) and $\pi$ is a sub-path of $\gamma$ such that $\pi\subset B$, then $\pi\in\mathbb{G}(B)$ (or $\pi\in\mathbb{G}_{*}(B)$). In addition, $\mathbb{G}(A)$ and $\mathbb{G}_{*}(A)$ are measurable with respect to edge weights inside $A$. Moreover, if $\gamma\in\mathbb{G}_{*}(u,v)$ and $\eta\setminus\gamma$ is open, then $\eta\in\mathbb{G}_{*}(u,v)$ as well.
For each $N\geq 1$, we define the family of $N$-annuli:
$\forall e\in{{\mathcal{E}}({\mathbb{Z}}^{d})},\quad{\mathrm{A}_{N}(e)}:=\Lambda_{3N}(e)\setminus\Lambda_{N}(e),\quad\partial{\mathrm{A}_{N}(e)}:=\partial\Lambda_{3N}(e)\cup\partial\Lambda_{N}(e).$ (2.1)
A path $\eta\subset{\mathrm{A}_{N}(e)}$ is called a crossing path of ${\mathrm{A}_{N}(e)}$ if it joints $\partial\Lambda_{N}(e)$ and $\partial\Lambda_{3N}(e)$. Let ${\mathscr{C}}({\mathrm{A}_{N}(e)})$ be the collection of all crossing paths of ${\mathrm{A}_{N}(e)}$. Suppose that $\gamma$ is a path crossing the annulus ${\mathrm{A}}_{N}(e)$ at least once ($\gamma$ is not necessarily inside ${\mathrm{A}}_{N}(e)$). Then, we denote by $\gamma^{{\bf i},e}\in{\mathscr{C}}({\mathrm{A}_{N}(e)})$ and $\gamma^{{\bf o},e}\in{\mathscr{C}}({\mathrm{A}_{N}(e)})$ the first and last sub-paths of $\gamma$ crossing ${\mathrm{A}}_{N}(e)$, respectively, ordered from ${\bf s}(\gamma)$ to ${\bf e}(\gamma)$ (see Figure 1-A).
(a) First and last crossing paths
(b) A short open path $\pi_{e}$ joining crossing paths
Figure 1. Illustration of crossing paths and the event $\mathcal{V}^{1}_{e}$
Fix $\rho$ and $c_{*}$ as in Lemmas 2.10 and 2.14, respectively, and define
$C_{*}:=1/c_{*}+(100\rho^{2})^{d};$ (2.2)
@@
$R_{e}:=\inf\{N\geq 1:\mathcal{V}_{N}^{1}(e)\cap\mathcal{V}_{N}^{2}(e)\text{ occurs}\}.$ (2.3)
Roughly speaking, $R_{e}$ is the radius that ensures the existence of an economically feasible bypass for the edge $e$ with a cost comparable to $R_{e}$ (see Figure 1-B). Additionally, as we will see below, $R_{e}$ has beneficial properties such as local dependence and a light-tailed distribution, using the robust connectivity of the supercritical percolation
@@
Given an edge $e$ in a modified geodesic, the following proposition allows us to build a detour that avoids $e$ with an economical cost that is comparable to the effective radius.
@@
$\eta_{e}:=\gamma_{x,z_{{\bf i}}}\cup\pi_{e}\cup\gamma_{z_{{\bf o}},y}.$
Therefore, $\eta_{e}\setminus\gamma=\pi_{e}$ is an open path. Together with $\gamma\in\mathbb{G}_{*}(x,y)$, we get $\eta_{e}\in\mathbb{G}_{*}(x,y)$, thanks to Remark 2.1. Notice that $|\eta_{e}\setminus\gamma|=|\pi_{e}|\leq C_{*}R_{e}$. Moreover, since $\gamma^{{\bf i},e}$ and $\gamma^{{\bf o},e}$ are the first and last crossing paths of ${\mathrm{A}}_{R_{e}}(e)$, one has $\gamma_{x,z_{\bf i}}\cap\Lambda_{R_{e}}(e)=\varnothing$ and $\gamma_{z_{\bf o},y}\cap\Lambda_{R_{e}}(e)=\varnothing$. In addition, $\pi_{e}\cap\Lambda_{R_{e}}(e)=\varnothing$ since $\pi_{e}\subset{\mathrm{A}}_{R_{e}}(e)$. Hence, $\eta_{e}\cap\Lambda_{R_{e}}(e)=\varnothing$ and so we obtain the claimed result. ∎
For any path $\gamma$, we denote by $\text{clo}(\gamma)$ the set of all closed edges of $\gamma$. The following result will help us gradually erase closed edges in a geodesic, turning it into an open path with a controllable cost. For more details, see in the proof of Proposition 5.1.
@@
Let $\gamma\in\mathbb{G}_{*}(x,y)$ with $x,y\in{\mathcal{C}}_{\infty}$ and let $e\in\text{clo}(\gamma)$ satisfying $\{x,y\}\not\subset\Lambda_{3R_{e}}(e)$. We consider three cases:
@@
Case 2: $x\in\Lambda_{3R_{e}}(e)$ and $y\notin\Lambda_{3R_{e}}(e)$. The path $\gamma$ crosses the annulus ${\mathrm{A}}_{R_{e}}(e)$ at least once. We call the last crossing path by $\gamma^{{\bf o},e}$. Using $\gamma\in\mathbb{G}_{*}$ and Remark 2.1, we get $\gamma^{{\bf o},e}\in{\mathscr{C}}({\mathrm{A}}_{R_{e}}(e))\cap\mathbb{G}_{*}(\Lambda_{C_{*}R_{e}}(e))$. As $x\in{\mathcal{C}}_{\infty}$, there exists an open path $\xi_{x,\infty}$ joining $x$ to $\infty$.
@@
$\eta_{e}:=\tilde{\xi}_{x,z_{\bf i}}\cup\tilde{\eta}_{e}\cup\gamma_{z_{\bf o},y}.$
Figure 2. Case 2a: $x\in\Lambda_{R_{e}}(e)$
We observe that the subpath $\eta_{e}\setminus\gamma\subset\{\tilde{\xi}_{x,z_{\bf i}}\cup\tilde{\eta}_{e}\}$ consists of only open edges, so $\eta_{e}\in\mathbb{G}_{*}(x,y)$. Moreover,
$|\eta_{e}\setminus\gamma|\leq|\tilde{\xi}_{x,z_{\bf i}}|+|\tilde{\eta}_{e}|\leq 2C_{*}R_{e},\quad\text{clo}(\eta_{e})\cap\Lambda_{R_{e}}(e)=\varnothing,$
since $\{\eta_{e}\cap\Lambda_{R_{e}}\}\subset\tilde{\xi}_{x,z_{\bf i}}$ is open. Hence, the result follows.
Case 2b: $x\in{\mathrm{A}}_{R_{e}}(e)$. Let $\gamma_{x,e}:=\gamma_{x,x_{e}}$ be the subpath of $\gamma$ joining $x$ to $e$. Let $\tilde{x}$ be the first point where $\xi_{x,\infty}$ touches $\partial\Lambda_{3R_{e}}(e)$, and so $\xi_{\tilde{x},x}\subset\Lambda_{3R_{e}}(e)$. Then, $\eta_{\tilde{x},e}:=\xi_{\tilde{x},x}\cup\gamma_{x,e}$ crosses the annulus ${\mathrm{A}}_{R_{e}}(e)$ at least once as $\xi_{\tilde{x},x}\cap\partial\Lambda_{3R_{e}}(e)\neq\varnothing$ and $\gamma_{x,e}\cap\partial\Lambda_{R_{e}}(e)\neq\varnothing$. Notice that since $\gamma\in\mathbb{G}_{*}$, we have $\gamma_{e,x}\in\mathbb{G}_{*}$, and so $\eta_{\tilde{x},e}\in\mathbb{G}_{*}$. Let $\gamma^{\textbf{i},e}\subset\eta_{\tilde{x},e}$ be the first crossing path of ${\mathrm{A}_{R_{e}}(e)}$. Thanks to Remark 2.1 again, $\gamma^{\textbf{i},e}\in{\mathscr{C}}({\mathrm{A}}_{R_{e}}(e))\cap\mathbb{G}_{*}(\Lambda_{C_{*}R_{e}}(e))$. By the definition of $\mathcal{V}^{1}_{R_{e}}(e)$, there exists a geodesic of ${\rm D}^{{\mathrm{A}}_{R_{e}}(e)}(\gamma^{{\bf i},e},\gamma^{{\bf o},e})$ inside ${\mathrm{A}}_{R_{e}}(e)$, denoted by $\tilde{\eta}_{e}$, that includes only open edges and satisfies $|\tilde{\eta}_{e}|\leq C_{*}R_{e}$. Suppose that $\tilde{\eta}_{e}$ intersects with $\gamma^{\textbf{i},e}$ and $\gamma^{\textbf{o},e}$ at $z_{\bf i}$ and $z_{\bf o}$, respectively.
@@
$\eta_{e}:=\gamma_{x,z_{\bf i}}\cup\tilde{\eta}_{e}\cup\gamma_{z_{\bf o},y}.$
(a) Subcase 2b: $z_{\bf i}\in\tilde{\eta}_{e}\cap\gamma_{x,e}$
(b) Subcase 2b: $z_{\bf i}\in\tilde{\eta}_{e}\cap\xi_{\tilde{x},x}$
Figure 3. Case 2b: $x\in{\mathrm{A}}_{R_{e}}(e)$
@@
$\eta_{e}:=\tilde{\xi}_{x,z_{\bf i}}\cup\tilde{\eta}_{e}\cup\gamma_{z_{\bf o},y}.$
Hence, $\tilde{\eta}_{e}\subset\eta_{e}\setminus\gamma\subset\tilde{\xi}_{x,z_{\bf i}}\cup\tilde{\eta}_{e}$ consists of only open edges, so $\eta_{e}\in\mathbb{G}_{*}(x,y)$. Also,
$|\eta_{e}\setminus\gamma|\leq|\tilde{\xi}_{x,z_{\bf i}}|+|\tilde{\eta}_{e}|\leq 2C_{*}R_{e},\quad\text{clo}(\eta_{e})\cap\Lambda_{R_{e}}(e)=\varnothing,$
as $\{\eta_{e}\cap\Lambda_{R_{e}}\}\subset\tilde{\xi}_{x,z_{\bf i}}$ is open. We complete the proof for this case.
Case 3: $x\notin\Lambda_{3R_{e}}(e)$ and $y\in\Lambda_{3R_{e}}(e)$. The proof is similar to that of Case 2 and is omitted. ∎
@@
$\forall\,e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}:\quad\hat{R}_{e}:=\min\{C_{*}R_{e},\,\log^{2}n\}.$ (2.4)
@@
$0\leq{\mathrm{T}}_{x,y}(\log^{2}n,t_{e^{c}})-{\mathrm{T}}_{x,y}(1,t_{e^{c}})\leq(\log^{2}n{\mathbb{I}}({\mathcal{U}}_{e})+\hat{R}_{e}){\mathbb{I}}(e\in{\gamma}),$ (2.5)
where ${\mathrm{T}}_{x,y}(t_{e},t_{e^{c}})$ denotes ${\mathrm{T}}(x,y)$ as a function of weights $t_{e}$ and $(t_{e^{\prime}})_{e^{\prime}\neq e}$, and $\gamma$ is a geodesic of ${\mathrm{T}}_{x,y}(1,t_{e^{c}})$, and

### v2

@@ 1.4. Organization and notation of this paper
  • Integer interval. Given an integer $t\geq 1$, we denote by $t]:=\{1,2,\ldots,t\}$.
@@
  • Box and its boundary. Let $x\in{\mathbb{Z}}^{d}$ and $t>0$, we will denote by $\Lambda_{t}(x):=x+[-t,t]^{d}\cap{\mathbb{Z}}^{d}$ the box centered at $x=(x_{1},\ldots,x_{d})\in{\mathbb{Z}}^{d}$ with radius $t$. For convenience, we briefly write $\Lambda_{t}=[-t,t]^{d}\cap{\mathbb{Z}}^{d}$ for $\Lambda_{t}(0)$. We define the boundary of $\Lambda_{t}(x)$ as $\partial\Lambda_{t}(x):=\Lambda_{t}(x)\setminus\Lambda_{t-1}(x)$.
  • Set of boxes. For $m\in{\mathbb{N}}$ and $X\subset{\mathbb{Z}}^{d}$, let ${\mathcal{B}}_{m}(X)$ denote the set of all boxes of side-length $m$ in $X$.
@@
  • ${\mathbb{Z}}^{d}$-path and set of ${\mathbb{Z}}^{d}$-paths. For any $\ell$, we say that a sequence $\gamma=(v_{0},\ldots,v_{\ell})$ is a ${\mathbb{Z}}^{d}$-path if for all $i\in[\ell],\|v_{i}-v_{i-1}\|_{1}=1$. The length of $\gamma$ is $\ell$, denoted by $|\gamma|$. For $1\leq i<j\leq\ell$, we denote by $\gamma_{v_{i},v_{j}}$ the subpath of $\gamma$ from $v_{i}$ to $v_{j}$. In addition, if $v_{i}\neq v_{j}$ for $i\neq j$, then we say that $\gamma$ is self-avoiding. From now on, we will shortly write a path in place of a self-avoiding ${\mathbb{Z}}^{d}$-path. For any path $\gamma$, we denote ${\bf s}(\gamma),{\bf e}(\gamma)$ the starting and ending vertices of $\gamma$, respectively. Given $U\subset{\mathbb{Z}}^{d}$, let $\mathcal{P}(U)$ be the set of all paths in $U$.
  • Paths concatenation. Given two paths $\gamma^{1},\gamma^{2}$ such that ${\bf e}(\gamma^{1})={\bf s}(\gamma^{2})$, we denote the paths concatenation of $\gamma^{1}$ and $\gamma^{2}$ by
    $\gamma^{1}\oplus\gamma^{2}=({\bf s}(\gamma^{1}),\ldots,{\bf e}(\gamma^{1})={\bf s}(\gamma^{2}),\ldots,{\bf e}(\gamma^{2})).$
@@
  • Open connection. Given $A,B,U\subset{\mathbb{Z}}^{d}$, we write $A\xleftrightarrow{U}B$ if there exists an open path inside $U$ connecting $A$ to $B$; otherwise, we write $A\not\xleftrightarrow{U}B$. When $U={\mathbb{Z}}^{d}$, we omit the symbol ${\mathbb{Z}}^{d}$ for simplicity.
@@
    and we thus denote diam(A)(A) the diameter of $A$ by
@@
  • Geodesic. Given $\{x,y\}\subset U\subset{\mathbb{Z}}^{d}$, we denote by $\gamma$ a geodesic between $x$ and $y$ of ${\mathrm{D}}^{U}(x,y)$ if $\gamma$ is an open path inside $U$ such that $|\gamma|={\mathrm{D}}^{U}(x,y)$. We also denote by $\eta$ a geodesic between $x$ and $y$ of ${\mathrm{T}}^{U}(x,y)$ if $\eta$ is a path inside $U$ such that ${\mathrm{T}}(\eta)={\mathrm{T}}^{U}(x,y)$. If there are several choices for $\gamma$ or $\eta$, we choose one of them according to a deterministic rule to break ties.
@@ 2. The effect of resampling
As mentioned in Introduction, the key to the concentration of the truncated passage time ${\mathrm{T}}_{n}$ is understanding the cost of resampling the edge weights along the geodesic of ${\mathrm{T}}_{n}$. To study this issue, given an edge $e$, we introduce the notion of effective radius $R_{e}$, which measures how large the change in passage time is when flipping the weight of $e$. The radius $R_{e}$ indeed guarantees an open bypass of $e$ whose length is comparable to $R_{e}$. Hence, it turns out that the cost of flipping the edge weight is given by (12). We then investigate the properties of the effective radii, including the local dependence and light-tailed decay distribution. These properties will play a central role in controlling the total cost of resampling edges.
@@ 2.1. Effective radius and its applications
$\mathbb{G}_{*}(u,v;A)$ $:=$ $\Big{\{}\gamma\in{\mathcal{P}}(A):{\bf s}(\gamma)=u,{\bf e}(\gamma)=v,\exists\,\pi\in\mathbb{G}(A):\gamma\setminus\pi\textrm{ is open}\Big{\}},$
$\mathbb{G}_{*}(A)$ $:=$ $\bigcup_{u,v\in A}\mathbb{G}_{*}(u,v;A),$
@@
The set of modified geodesics $\mathbb{G}_{*}$ arises from our goal of controlling the difference between the graph distance and the first truncated passage time. To achieve this, we build an inductive process that, step by step, replaces certain sub-geodesics of the first truncated passage time with open paths (see Proposition 5.1 for more details). Consequently, at each step, the updated geodesic is no longer in $\mathbb{G}$, but instead belongs to $\mathbb{G}_{*}$.
Remark 2.2.
Given $B\subset A\subset{\mathbb{Z}}^{d}$, if $\gamma\in\mathbb{G}(A)$ (or $\gamma\in\mathbb{G}_{*}(A)$) and $\pi$ is a subpath of $\gamma$ such that $\pi\subset B$, then $\pi\in\mathbb{G}(B)$ (or $\pi\in\mathbb{G}_{*}(B)$). In addition, $\mathbb{G}(A)$ and $\mathbb{G}_{*}(A)$ are measurable with respect to edge weights inside $A$. Moreover, if $\gamma\in\mathbb{G}_{*}(u,v)$ and $\eta$ is a path from $u$ to $v$ such that $\eta\setminus\gamma$ is open, then $\eta\in\mathbb{G}_{*}(u,v)$ as well.
Recall that for each $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$, we fix a deterministic rule to write $e=(x_{e},y_{e})$ so that $\|x_{e}\|_{1}<\|y_{e}\|_{1}$. For each $N\geq 1$ and $e=(x_{e},y_{e})$, we define $\Lambda_{N}(e):=\Lambda_{N}(x_{e})$ and the family of $N$-annuli:
$\forall e\in{{\mathcal{E}}({\mathbb{Z}}^{d})},\quad{\mathrm{A}_{N}(e)}:=\Lambda_{3N}(e)\setminus\Lambda_{N}(e),\quad\partial{\mathrm{A}_{N}(e)}:=\partial\Lambda_{3N}(e)\cup\partial\Lambda_{N}(e).$ (17)
A path $\eta\subset{\mathrm{A}_{N}(e)}$ is called a crossing path of $\mathrm{A}_{N}(e)$} if it joints $\partial\Lambda_{N}(e)$ and $\partial\Lambda_{3N}(e)$. Let ${\mathscr{C}}({\mathrm{A}_{N}(e)})$ be the collection of all crossing paths of $\mathrm{A}_{N}(e)$}. Suppose that $\gamma$ is a path crossing the annulus ${\mathrm{A}}_{N}(e)$ at least once ($\gamma$ is not necessarily inside ${\mathrm{A}}_{N}(e)$). Then, we denote by $\gamma^{{\bf i},e}\in{\mathscr{C}}({\mathrm{A}_{N}(e)})$ and $\gamma^{{\bf o},e}\in{\mathscr{C}}({\mathrm{A}_{N}(e)})$ the first and last subpaths of $\gamma$ crossing ${\mathrm{A}}_{N}(e)$, respectively, ordered from ${\bf s}(\gamma)$ to ${\bf e}(\gamma)$ (see Figure 1-A).
(a) First and last crossing paths.
(b) A short open path $\pi_{e}$ joining crossing paths.
Figure 1. Illustration of crossing paths and the event $\mathcal{V}^{1}_{e}$.
Fix $\rho$ and $c_{*}$ as in Lemmas 2.12 and 2.17, respectively, and define
$C_{*}:=1/c_{*}+(100\rho^{2})^{d};$ (18)
@@
$R_{e}:=\inf\{N\geq 1:\mathcal{V}_{N}^{1}(e)\cap\mathcal{V}_{N}^{2}(e)\text{ occurs}\}.$ (19)
Roughly speaking, $R_{e}$ is the radius that ensures the existence of an economically feasible bypass for any path in $\mathbb{G}_{*}$ crossing the annulus with a cost comparable to $R_{e}$ (see Figure 1-B). Additionally, as we will see below, $R_{e}$ has beneficial properties such as local dependence and a light-tailed distribution, using the robust connectivity of the supercritical percolation.
Remark 2.4.
In [9], Nakajima and we also introduced a simplified version of the effective radius to study the Lipschitz continuity of the time constant, motivated by an earlier draft of this paper. Here, we highlight the key difference between these notions. In the present paper, the event ${\mathcal{V}}^{1}_{N}$ involves the graph distance between two more complex paths, rather than focusing solely on geodesics as in [9]. This modification, along with the introduction of ${\mathcal{V}}^{2}_{N}$, aims to better control the discrepancy between the graph distance and the truncated first-passage time. Notably, this change allows us to obtain exponentially large deviation estimates for the effective radius, rather than the polynomial bounds established in [9]. We emphasize that while the moment bounds for this discrepancy were sufficient in [9], they are no longer adequate in the present paper.
@@
We can now state the large deviation estimate for these effective radii:
@@
The proof of Proposition 2.6 is provided in Section 2.2. Given an edge $e$ in a modified geodesic, the following proposition allows us to build a detour that avoids $e$ with an economical cost that is comparable to the effective radius.
@@
$\eta_{e}:=\gamma_{x,z_{{\bf i}}}\oplus\pi_{e}\oplus\gamma_{z_{{\bf o}},y},$
where we recall that, for two paths $\sigma^{1}$ and $\sigma^{2}$ with ${\bf e}(\sigma^{1})={\bf s}(\sigma^{2})$, the operator $\sigma^{1}\oplus\sigma^{2}$ denotes their concatenation. Therefore, $\eta_{e}\setminus\gamma=\pi_{e}$ is an open path. Together with $\gamma\in\mathbb{G}_{*}(x,y)$, we get $\eta_{e}\in\mathbb{G}_{*}(x,y)$, thanks to Remark 2.2. Notice that $|\eta_{e}\setminus\gamma|=|\pi_{e}|\leq C_{*}R_{e}$. Moreover, since $\gamma^{{\bf i},e}$ and $\gamma^{{\bf o},e}$ are the first and last crossing paths of ${\mathrm{A}}_{R_{e}}(e)$, one has $\gamma_{x,z_{\bf i}}\cap\Lambda_{R_{e}}(e)=\varnothing$ and $\gamma_{z_{\bf o},y}\cap\Lambda_{R_{e}}(e)=\varnothing$. In addition, $\pi_{e}\cap\Lambda_{R_{e}}(e)=\varnothing$ since $\pi_{e}\subset{\mathrm{A}}_{R_{e}}(e)$. Hence, $\eta_{e}\cap\Lambda_{R_{e}}(e)=\varnothing$ and so we obtain the claimed result. ∎
For any path $\gamma$, we denote by ${\mathrm{clo}}(\gamma)$ the set of all closed edges of $\gamma$. The following result will help us gradually erase closed edges in a geodesic, turning it into an open path with a controllable cost. For more details, see in the proof of Proposition 5.1.
@@
Let $\gamma\in\mathbb{G}_{*}(x,y)$ with $x,y\in{\mathcal{C}}_{\infty}$ and let $e\in{\mathrm{clo}}(\gamma)$ satisfying $\{x,y\}\not\subset\Lambda_{3R_{e}}(e)$. We consider two cases:
@@
Case 2: there is only $x$ or $y$ in $\Lambda_{3R_{e}}(e)$. By the symmetry, we suppose that $x\in\Lambda_{3R_{e}}(e)$ and $y\notin\Lambda_{3R_{e}}(e)$. The path $\gamma$ crosses the annulus ${\mathrm{A}}_{R_{e}}(e)$ at least once. We call the last crossing path by $\gamma^{{\bf o},e}$. Using $\gamma\in\mathbb{G}_{*}$ and Remark 2.2, we get $\gamma^{{\bf o},e}\in{\mathscr{C}}({\mathrm{A}}_{R_{e}}(e))\cap\mathbb{G}_{*}(\Lambda_{C_{*}R_{e}}(e))$. As $x\in{\mathcal{C}}_{\infty}$, there exists an open path $\xi_{x,\infty}$ joining $x$ to $\infty$.
@@
$\eta_{e}:=\tilde{\xi}_{x,z_{\bf i}}\oplus\tilde{\eta}_{e}\oplus\gamma_{z_{\bf o},y}.$
Figure 2. Construction of the detour $\eta_{e}$ avoiding the closed edge $e$, obtained by concatenating the paths $\tilde{\xi}_{x,z_{\bf i}}$, $\tilde{\eta}_{e}$, and $\gamma_{z_{\bf o},y}$ when $x\in\Lambda_{R_{e}}(e)$.
We observe that the subpath $\eta_{e}\setminus\gamma\subset\{\tilde{\xi}_{x,z_{\bf i}}\oplus\tilde{\eta}_{e}\}$ consists of only open edges, so $\eta_{e}\in\mathbb{G}_{*}(x,y)$. Moreover,
$|\eta_{e}\setminus\gamma|\leq|\tilde{\xi}_{x,z_{\bf i}}|+|\tilde{\eta}_{e}|\leq 2C_{*}R_{e},\quad{\mathrm{clo}}(\eta_{e})\cap\Lambda_{R_{e}}(e)=\varnothing,$
since $\{\eta_{e}\cap\Lambda_{R_{e}}(e)\}\subset\tilde{\xi}_{x,z_{\bf i}}$ is open. Hence, the result follows.
Case 2b: $x\in{\mathrm{A}}_{R_{e}}(e)$. Let $\gamma_{x,e}:=\gamma_{x,x_{e}}$ be the subpath of $\gamma$ joining $x$ to $e$. Let $\tilde{x}$ be the first point where $\xi_{x,\infty}$ touches $\partial\Lambda_{3R_{e}}(e)$, and so $\xi_{\tilde{x},x}\subset\Lambda_{3R_{e}}(e)$. Then, $\eta_{\tilde{x},e}:=\xi_{\tilde{x},x}\oplus\gamma_{x,e}$ crosses the annulus ${\mathrm{A}}_{R_{e}}(e)$ at least once as $\xi_{\tilde{x},x}\cap\partial\Lambda_{3R_{e}}(e)\neq\varnothing$ and $\gamma_{x,e}\cap\partial\Lambda_{R_{e}}(e)\neq\varnothing$. Notice that since $\gamma\in\mathbb{G}_{*}$, we have $\gamma_{e,x}\in\mathbb{G}_{*}$, and so $\eta_{\tilde{x},e}\in\mathbb{G}_{*}$. Let $\gamma^{\textbf{i},e}\subset\eta_{\tilde{x},e}$ be the first crossing path of $\mathrm{A}_{R_{e}}(e)$}. Thanks to Remark 2.2 again, $\gamma^{\textbf{i},e}\in{\mathscr{C}}({\mathrm{A}}_{R_{e}}(e))\cap\mathbb{G}_{*}(\Lambda_{C_{*}R_{e}}(e))$. By the definition of $\mathcal{V}^{1}_{R_{e}}(e)$, there exists a geodesic of ${\rm D}^{{\mathrm{A}}_{R_{e}}(e)}(\gamma^{{\bf i},e},\gamma^{{\bf o},e})$ inside ${\mathrm{A}}_{R_{e}}(e)$, denoted by $\tilde{\eta}_{e}$, that includes only open edges and satisfies $|\tilde{\eta}_{e}|\leq C_{*}R_{e}$. Suppose that $\tilde{\eta}_{e}$ intersects with $\gamma^{\textbf{i},e}$ and $\gamma^{\textbf{o},e}$ at $z_{\bf i}$ and $z_{\bf o}$, respectively.
@@
$\eta_{e}:=\gamma_{x,z_{\bf i}}\oplus\tilde{\eta}_{e}\oplus\gamma_{z_{\bf o},y}.$
(a) if $z_{\bf i}\in\tilde{\eta}_{e}\cap\gamma_{x,e}$, then $\eta_{e}$ is obtained by concatenating $\gamma_{x,z_{\bf i}}$, $\tilde{\eta}_{e}$, and $\gamma_{z_{\bf o},y}$.
(b) if $z_{\bf i}\in\tilde{\eta}_{e}\cap\xi_{\tilde{x},x}$, then $\eta_{e}$ is obtained by concatenating $\tilde{\xi}_{x,z_{\bf i}},\tilde{\eta}_{e}$, and $\gamma_{z_{\bf o},y}$.
Figure 3. Construction of the detour $\eta_{e}$ avoiding the closed edge $e$ when $x\in{\mathrm{A}}_{R_{e}}(e)$.
@@
$\eta_{e}:=\tilde{\xi}_{x,z_{\bf i}}\oplus\tilde{\eta}_{e}\oplus\gamma_{z_{\bf o},y}.$
Hence, $\tilde{\eta}_{e}\subset\eta_{e}\setminus\gamma\subset\{\tilde{\xi}_{x,z_{\bf i}}\oplus\tilde{\eta}_{e}\}$ consists of only open edges, so $\eta_{e}\in\mathbb{G}_{*}(x,y)$. Also,
$|\eta_{e}\setminus\gamma|\leq|\tilde{\xi}_{x,z_{\bf i}}|+|\tilde{\eta}_{e}|\leq 2C_{*}R_{e},\quad{\mathrm{clo}}(\eta_{e})\cap\Lambda_{R_{e}}(e)=\varnothing,$
as $\{\eta_{e}\cap\Lambda_{R_{e}}(e)\}\subset\tilde{\xi}_{x,z_{\bf i}}$ is open. We complete the proof for this case.
@@
$\forall\,e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}:\quad\hat{R}_{e}:=\min\{C_{*}R_{e},\,\log^{2}n\}.$ (20)
@@
$0\leq({\mathrm{T}}_{x,y}(\log^{2}n,t_{e^{c}})-{\mathrm{T}}_{x,y}(1,t_{e^{c}})){\mathbb{I}}(t_{e}=1)\leq(\log^{2}n{\mathbb{I}}({\mathcal{U}}_{e})+\hat{R}_{e}){\mathbb{I}}(e\in{\gamma}),$ (21)
where ${\mathrm{T}}_{x,y}(t_{e},t_{e^{c}})$ denotes ${\mathrm{T}}(x,y)$ as a function of weights $t_{e}$ and $(t_{e^{\prime}})_{e^{\prime}\neq e}$, and $\gamma$ is a geodesic of ${\mathrm{T}}_{x,y}(t_{e},t_{e^{c}})$, and